OpenStudy (anonymous):

How would I solve for U? Where U is a real number. So confused.... Help? √u+21=5

OpenStudy (anonymous):

which is the problem: \(\large \sqrt u +21=5 \) or \(\large \sqrt{u+21}=5 \) ???

OpenStudy (agent0smith):

\[\large \sqrt{u+21}=5\] I'd guess this one since the other gives sqrtu = -16, and U is real. Try squaring both sides to "undo" the square root: \[\large \sqrt{u+21}^2=5^2\]

OpenStudy (anonymous):

I'm sorry I don't understand :|

OpenStudy (anonymous):

how would squaring both sides undo the root?

OpenStudy (agent0smith):

Because squaring is like the opposite of square root, same way dividing is the opposite of multiplying, subtraction is the opposite of addition.

OpenStudy (agent0smith):

Or if you've done some exponents, \[\LARGE \sqrt{u+21}^2 = \left( (u+21)^{\frac{ 1 }{ 2 } }\right) ^2 = (u+21)^{\frac{ 2 }{2} }\]